# Shrinking the Generation-Verification Gap with Weak Verifiers

### v1

@@ Abstract
Verifiers can improve language model (LM) capabilities by scoring and ranking responses from a pool of generated candidates. Currently, high-quality verifiers are either unscalable (e.g., humans) or limited in utility (e.g., tools like Lean for formal proofs). While LM judges and reward models have become broadly useful as general-purpose verifiers, a significant performance gap remains between them and oracle verifiers (i.e. verifiers with perfect accuracy). To help close this gap, we introduce Weaver, a framework for designing a strong verifier by combining multiple weak, imperfect verifiers. First we find that weighted ensembles of verifiers, which typically require learning from labeled data, significantly outperform unweighted combinations due to differences in verifier accuracies. To reduce the dependency on labeled data, Weaver leverages weak supervision to estimate each verifier’s accuracy and combines their outputs into a unified score that better reflects true response quality. However, directly applying weak supervision algorithms poses several challenges, including inconsistent verifier output formats and handling low-quality verifiers. Weaver addresses these challenges by using dataset statistics to normalize outputs and filter specific verifiers. We study the effectiveness of Weaver in test-time repeated sampling settings, where a model generates multiple candidate responses and selects one from among them. Our evaluations demonstrate that Weaver significantly improves over $Pass@1$—the performance when simply selecting the first candidate response—across several reasoning and math tasks, achieving o3-mini-level accuracy with Llama 3.3 70B Instruct (a much cheaper non-reasoning model) as the generator, and an ensemble of 70B or smaller judge and reward models as the verifiers (87.7% average). This gain mirrors the jump achieved between GPT-4o and o3-mini (69.0% vs. 86.7%), which required extensive finetuning and post-training interventions. To reduce the computational costs of running verifier ensembles for Weaver, we train a compact 400M cross-encoder using Weaver’s combined output scores. This distilled model retains 98.7% of Weaver’s full accuracy while reducing verification compute by up to 99.97%.
@@ 1 Introduction
    Verification is expensive to deploy at inference. Verification can dominate inference-time costs (singhi2025solveverifycomputeoptimalproblem, ; liu20251bllmsurpass405b, ), since each verifier must process both the problem and its candidate response(s) lightman2023letsverifystepstep , often evaluating intermediate steps lightman2023letsverifystepstep and multiple solution paths snell2024scalingllmtesttimecompute . In fact, achieving gains over unverified generation (i.e. majority voting) can require 10×10$\times$10 × to 128×128$\times$128 × the inference compute per query (singhi2025whentosolve, ; lifshitz2025multiagentverificationscalingtesttime, ; zhao2025sample, ; chen2025sets, ).
@@ 4.1 How to aggregate multiple verifiers: weighted vs unweighted ensembles
An alternative is to use a weighted ensemble. One approach is to use a labeled dataset to identify and use the top-performing verifier, effectively assigning a weight of 00 to discarded verifiers. Other strategies include using Logistic Regression or a Naive Bayes classifier, where the scoring function $f(q_{i},r_{ij})$ is the probability $\Pr(y_{ij}=1|s_{ij1},\dots,s_{ijm})$. These classifiers are fit using labeled data and can be either modeled as a logistic function or factorized using Bayes’ rule and independence assumptions, respectively.
@@ 4.2.1 Weak Supervision Algorithm
This is a consistency property that holds regardless of the conditional independence assumption, and we can write this equation for each of the $m$ $S_{i}$’s. Because we know that the accuracy parameters should follow equations 2 and 3, we can construct an objective function that aims to minimize the difference between the left and right hand sides of these equations. We write this efficiently in matrix notation. Let $P\in\mathbb{R}^{2\times 2}$ be a diagonal matrix with diagonal $[\Pr(Y=0)\;\Pr(Y=1)]$. Define $\mu\in\mathbb{R}^{m\times 2}$ to be the matrix of accuracy parameters, and define $O\in\mathbb{R}^{2m\times 2m}$ to be a matrix over the joint probabilities of pairs of $S_{i},S_{j}$; more formally:
@@
We optimize 5 using gradient descent to estimate the verifier accuracy parameters. These estimates are then used in Eq. 1 to select the response with the highest estimated posterior. To further improve modeling of verifier accuracies, we explore whether partitioning the query distribution by empirical difficulty can yield better weak supervision estimates. As detailed in Appendix B.4, we cluster queries based on the observed ratio of correct to incorrect generations, and fit a separate Weaver model within each difficulty bucket. We provide more details in Sec. B.1.
@@ 5 Results
Datasets, Verifiers, and Baselines    Our reward models range in size from 8B to 72B, are all open-source, and are obtained from RewardBench lambert2024rewardbenchevaluatingrewardmodels , a popular evaluation tool for reward models. We prompt open-source language models from Chatbot Arena (chiang2024chatbot, ) to serve as judges. Unless specified, we use Llama 3.3 70B Instruct to generate responses and use all 33 reward models and judges. We evaluate on MATH500, GPQA Diamond, MMLU College, and MMLU Pro. See Sec. C.1 for more details.
@@ 5.1 Weaver Shrinks the Gap with Frontier LMs
In Table 1, we evaluate Weaver along with baseline verification methods, the first sample performance of frontier LMs, and the Pass@100 metric. We use LlaMA 3.3 70B Instruct to generate $K=100$ responses per query. We find that Weaver’s weighted ensembling of multiple verifiers allows us to outperform majority vote by $15.5\%$ and come within $4.2\%$ of the Pass@100 oracle metric. Furthermore, Weaver rivals the performance of frontier reasoning models—coming within $0.5\%$ of OpenAI’s o3-mini (OpenAIo3mini2025, )—even though we use a non-reasoning model for generation.
@@
Methodology Generations ($K$) Datasets Average
@@
Baselines First Sample 1 78.0% 42.9% 82.6% 69.9% 68.4%
@@
Frontier Approaches GPT-4o (openaigpt4, ) 1 77.4% 35.9% 87.1% 75.4% 69.0%
@@ 5.2 Weaver Improves Compute-Accuracy Trade-Off for Scaling
By proposing to combine multiple weak verifiers instead of one, we introduce yet another axis for test-time scaling. In this section, we study how well scaling verification with Weaver interacts with common previously studied axes for verification, summarized in Table 2.
@@
(1) Scaling Candidate Generations: we study the performance of verification methods as we increase the number of repeated samples in Fig. 3. Based on prior work bradley1952rank ; chen2021evaluating , as the number of responses increases, we are more likely to see a correct response (i.e. Pass@K increases), and hence more likely to select a correct response given a good verification strategy. However, differences in verification translate into different scaling rates. We evaluate the performance of Weaver and baselines for $K=2^{0}$ to $2^{10}$, comparing to o3-mini and Pass@K as well. Across all tasks, Weaver yields the most substantial gains when scaling the number of generations. Weaver consistently narrows the generation-verification gap with the oracle upper bound (Pass@K) while alternative verification strategies plateau after a few generations. The effect is particularly pronounced on difficult tasks like GPQA. We detail the scaling trends observed in Fig. 3 in Sec. C.3.
@@
In Table 3, we study how Weaver applied on smaller models (both verifiers and for generating responses) can allow us to match the performance of larger models, enabling weak-to-strong verification. We consider an 8B setting—using LlaMA 3.1 8B to generate responses along with 8B verifiers—and compare this to a 70B setting (LlaMA 3.3 70B Instruct, 8B-72B verifiers) as well as o3-mini. We see that Weaver applied at the 8B scale comes within $1.6\%$ of the majority vote baseline at the 70B scale, and Weaver at 70B surpasses o3-mini by $1.0\%$, demonstrating a weak-to-strong verification phenomenon. Verifier calibration details are available in Appendix C.5.
@@
Two axes for scaling verification are (1) the number of verifiers used and (2) the number of scores sampled from each verifier. Fig. 4 shows how performance changes as we ensemble 1 to 15 verifiers using both naive averaging and Weaver. Verifiers are greedily added in order of individual accuracy, from highest to lowest. Aggregating more verifiers improves performance by up to $8.5\%$ over the top-1 verifier. As shown in Fig. 4, Weaver consistently outperforms naive ensemble averaging across both Oracle Top-5 Verifiers and Total Verifiers configurations for verifier ensembling, with improvements ranging from +2.4% to +10.1% across all datasets. The performance gains are particularly pronounced on GPQA Diamond (+10.1%) and MMLU Pro (+5.1%), demonstrating Weaver’s effectiveness in aggregating verifier signals through learned weights rather than simple averaging. However, gains diminish as more models are added—reflecting the classic ensemble bias-variance tradeoff: initial improvements stem from variance reduction, while additional verifiers contribute redundant signal due to correlated biases on hard examples abebuchanan2024 . We compare alternative score calibration strategies beyond Weaver’s binary transformation in Appendix C.7, and find that the default binarization yields the strongest downstream selection performance. We also explore scaling the number of scores per verifier—via prompt tuning or temperature variation—in Appendix C.5. While this yields modest improvements, increasing verifier count remains the more effective strategy. That said, both methods are complementary and can be combined for further gains.
@@
[…]
@@
We study how performance scales in the total compute used for both verification and repeated generations. Figure 5 shows the relationship between inference-time compute and success rate for different generation-verification systems. For each method, we scale the number of generations exponentially from 1 to 100 and plot the required inference compute for generation and verification together versus the success rate. Note that Fig. 5 differs from Fig. 3, since Majority Voting requires 00 verification inference calls while Weaver requires 30+ calls for the weak verifiers. We find that Weaver achieves the highest maximum success rate; notably, majority voting plateaus at around $2^{2}$ to $2^{3}$ ExaFLOPs per query while Weaver continues scaling until 512 ExaFLOPs. However, the additional compute required for Weaver can be prohibitive. We explore how to reduce this computational burden while retaining Weaver’s performance in the next section.
@@ 6 Weaver Distillation: Improving Verification Efficiency at Inference
Figure 6 shows the performance of Weaver on the Llama-70B generations against the cross-encoder on GPQA Diamond. Across tasks, we find that the distilled cross-encoder is able to capture 98.2% of the performance of Weaver. When running Weaver with all the verifiers, it costs 35.35 exaFLOPs for each query’s set of 100 samples. Running a 400M cross-encoder costs 1.01 exaFLOPs for evaluating 100 samples and reduces compute cost by more than three orders of magnitude, saving 99.97% of the FLOPs originally required for running the 70B verifiers. We also outperform majority voting by 23.2% while only incurring a 0.57% increased inference cost over only generating the responses. We see similar results for additional datasets in Figure 22 (Sec. C.6).
@@ B.1 Weak Supervision Model
$\Pr(S_{k}=\bar{s}_{ijk}\mid Y=y_{ij})=\begin{cases}w_{k,1},&\text{if }y_{ij}=1\text{ and }\bar{s}_{ijk}=1,\\ 1-w_{k,1},&\text{if }y_{ij}=1\text{ and }\bar{s}_{ijk}=0,\\ w_{k,0},&\text{if }y_{ij}=0\text{ and }\bar{s}_{ijk}=0,\\ 1-w_{k,0},&\text{if }y_{ij}=0\text{ and }\bar{s}_{ijk}=1.\end{cases}$
@@
Eq. 6 requires evaluating the full conditional likelihood:
@@
which is a joint distribution over $m$ binary random variables. Since each verifier $S_{k}\in\{0,1\}$ is binary, then there are $2^{m}$ possible verifier output configurations for $S\in\{0,1\}^{m}$. This results in $12^{m}-$ 1 free parameters per class label to construct the distribution $\Pr(S=\bar{s}_{ij}\mid y_{ij})$.
@@
Then, Eq. 6 simplifies to a Naive Bayes-style estimator:
@@ B.2 Adapting Weak Supervision to the Verification Setting
Given a set of verifiers, we elaborate on the design choices behind the weak supervision model described in Sec. 4. In particular, we describe challenges around normalization, binarization and filtering out low-quality verifiers. In Section B.3, we describe Weaver’s approach to normalizing, binarizing, and filtering verifiers.
@@ B.2.2 Binarization
The weak supervision algorithm described in the prior section requires binary verifier outputs. This is naturally suited for judge-style verifiers—such as language models prompted to answer yes/no questions—which output discrete ${0,1}$ labels. However, many verifiers, especially reward models, emit continuous scores and often vary in scale and calibration. This raises a key design question: should we input these scores to the weak supervision model as-is, or should we binarize them?
Using continuous scores retains fine-grained information about the confidence of each verifier. This can improve ranking-based performance metrics such as AUC and may allow the weak supervision model to better resolve disagreements among verifiers. However, it introduces challenges when combining signals across verifiers with inconsistent calibration or scale: a score of 0.8 may have different meanings for different verifiers. As seen in Figure Fig. 7 different verifiers exhibit different score distributions even when evaluating the same set of responses. Some verifiers are sharply bimodal, others skew heavily toward low or high scores, and some produce nearly flat or noisy distributions.
@@
Figure 8: Same conventions as Fig. 9 where the performance reported is the area under the curve of the ROC.

### v2

@@ Abstract
Verifiers can improve language model (LM) capabilities by scoring and ranking responses from a pool of generated candidates. Currently, high-quality verifiers are either unscalable (e.g., humans) or limited in utility (e.g., tools like Lean for formal proofs). While LM judges and reward models have become broadly useful as general-purpose verifiers, a significant performance gap remains between them and oracle verifiers (i.e. verifiers with perfect accuracy). To help close this gap, we introduce Weaver, a framework for designing a strong verifier by combining multiple weak, imperfect verifiers. First we find that weighted ensembles of verifiers, which typically require learning from labeled data, significantly outperform unweighted combinations due to differences in verifier accuracies. To reduce the dependency on labeled data, Weaver leverages weak supervision to estimate each verifier’s accuracy and combines their outputs into a unified score that better reflects true response quality. However, directly applying weak supervision algorithms poses several challenges, including inconsistent verifier output formats and handling low-quality verifiers. Weaver addresses these challenges by using dataset statistics to normalize outputs and filter specific verifiers. We study the effectiveness of Weaver in test-time repeated sampling settings, where a model generates multiple candidate responses and selects one from among them. Our evaluations demonstrate that Weaver significantly improves over Pass@1Pass@1—the performance when simply selecting the first candidate response—across several reasoning and math tasks, achieving o3-mini-level accuracy with Llama 3.3 70B Instruct (a much cheaper non-reasoning model) as the generator, and an ensemble of 70B or smaller judge and reward models as the verifiers (87.7% average). This gain mirrors the jump achieved between GPT-4o and o3-mini (69.0% vs. 86.7%), which required extensive finetuning and post-training interventions. To reduce the computational costs of running verifier ensembles for Weaver, we train a compact 400M cross-encoder using Weaver’s combined output scores. This distilled model retains 98.7% of Weaver’s full accuracy while reducing verification compute by up to 99.97%.
@@ 1 Introduction
    Verification is expensive to deploy at inference. Verification can dominate inference-time costs (singhi2025solveverifycomputeoptimalproblem, ; liu20251bllmsurpass405b, ), since each verifier must process both the problem and its candidate response(s) lightman2023letsverifystepstep , often evaluating intermediate steps lightman2023letsverifystepstep and multiple solution paths snell2024scalingllmtesttimecompute . In fact, achieving gains over unverified generation (i.e. majority voting) can require $10\times$ to $128\times$ the inference compute per query (singhi2025whentosolve, ; lifshitz2025multiagentverificationscalingtesttime, ; zhao2025sample, ; chen2025sets, ).
@@ 4.1 How to aggregate multiple verifiers: weighted vs unweighted ensembles
An alternative is to use a weighted ensemble. One approach is to use a labeled dataset to identify and use the top-performing verifier, effectively assigning a weight of 0 to discarded verifiers. Other strategies include using Logistic Regression or a Naive Bayes classifier, where the scoring function $f(q_{i},r_{ij})$ is the probability $\Pr(y_{ij}=1|s_{ij1},\dots,s_{ijm})$. These classifiers are fit using labeled data and can be either modeled as a logistic function or factorized using Bayes’ rule and independence assumptions, respectively.
@@ 4.2.1 Weak Supervision Algorithm
This is a consistency property that holds regardless of the conditional independence assumption, and we can write this equation for each of the $m$ $S_{i}$’s. Because we know that the accuracy parameters should follow equations 2 and 3, we can construct an objective function that aims to minimize the difference between the left and right hand sides of these equations. We write this efficiently in matrix notation. Let $P\in\mathbb{R}^{2\times 2}$ be a diagonal matrix with diagonal $\Pr(Y=0)\;\Pr(Y=1)]$. Define $\mu\in\mathbb{R}^{m\times 2}$ to be the matrix of accuracy parameters, and define $O\in\mathbb{R}^{2m\times 2m}$ to be a matrix over the joint probabilities of pairs of $S_{i},S_{j}$; more formally:
@@
We optimize 5 using gradient descent to estimate the verifier accuracy parameters. These estimates are then used in Eq.˜1 to select the response with the highest estimated posterior. To further improve modeling of verifier accuracies, we explore whether partitioning the query distribution by empirical difficulty can yield better weak supervision estimates. As detailed in Appendix B.4, we cluster queries based on the observed ratio of correct to incorrect generations, and fit a separate Weaver model within each difficulty bucket. We provide more details in Sec.˜B.1.
@@ 5 Results
Datasets, Verifiers, and Baselines    Our reward models range in size from 8B to 72B, are all open-source, and are obtained from RewardBench lambert2024rewardbenchevaluatingrewardmodels , a popular evaluation tool for reward models. We prompt open-source language models from Chatbot Arena (chiang2024chatbot, ) to serve as judges. Unless specified, we use Llama 3.3 70B Instruct to generate responses and use all 33 reward models and judges. We evaluate on MATH500, GPQA Diamond, MMLU College, and MMLU Pro. See Sec.˜C.1 for more details.
@@ 5.1 Weaver Shrinks the Gap with Frontier LMs
In Table˜1, we evaluate Weaver along with baseline verification methods, the first sample performance of frontier LMs, and the Pass@100 metric. We use LlaMA 3.3 70B Instruct to generate $K=100$ responses per query. We find that Weaver’s weighted ensembling of multiple verifiers allows us to outperform majority vote by $15.5\%$ and come within $4.2\%$ of the Pass@100 oracle metric. Furthermore, Weaver rivals the performance of frontier reasoning models—coming within $0.5\%$ of OpenAI’s o3-mini (OpenAIo3mini2025, )—even though we use a non-reasoning model for generation.
@@
Methodology
Generations ($K$)
Datasets
Average
@@
Baselines
First Sample 1 78.0% 42.9% 82.6% 69.9% 68.4%
@@
Frontier
Approaches
GPT-4o (openaigpt4, ) 1 77.4% 35.9% 87.1% 75.4% 69.0%
@@ 5.2 Weaver Improves Compute-Accuracy Trade-Off for Scaling
By proposing to combine multiple weak verifiers instead of one, we introduce yet another axis for test-time scaling. In this section, we study how well scaling verification with Weaver interacts with common previously studied axes for verification, summarized in Table˜2.
@@
(1) Scaling Candidate Generations: we study the performance of verification methods as we increase the number of repeated samples in Fig.˜3. Based on prior work bradley1952rank ; chen2021evaluating , as the number of responses increases, we are more likely to see a correct response (i.e. Pass@K increases), and hence more likely to select a correct response given a good verification strategy. However, differences in verification translate into different scaling rates. We evaluate the performance of Weaver and baselines for $K=2^{0}$ to $2^{10}$, comparing to o3-mini and Pass@K as well. Across all tasks, Weaver yields the most substantial gains when scaling the number of generations. Weaver consistently narrows the generation-verification gap with the oracle upper bound (Pass@K) while alternative verification strategies plateau after a few generations. The effect is particularly pronounced on difficult tasks like GPQA. We detail the scaling trends observed in Fig.˜3 in Sec.˜C.3.
@@
In Table˜3, we study how Weaver applied on smaller models (both verifiers and for generating responses) can allow us to match the performance of larger models, enabling weak-to-strong verification. We consider an 8B setting—using LlaMA 3.1 8B to generate responses along with 8B verifiers—and compare this to a 70B setting (LlaMA 3.3 70B Instruct, 8B-72B verifiers) as well as o3-mini. We see that Weaver applied at the 8B scale comes within $1.6\%$ of the majority vote baseline at the 70B scale, and Weaver at 70B surpasses o3-mini by $1.0\%$, demonstrating a weak-to-strong verification phenomenon. Verifier calibration details are available in Appendix C.5.
@@
Two axes for scaling verification are (1) the number of verifiers used and (2) the number of scores sampled from each verifier. Fig.˜4 shows how performance changes as we ensemble 1 to 15 verifiers using both naive averaging and Weaver. Verifiers are greedily added in order of individual accuracy, from highest to lowest. Aggregating more verifiers improves performance by up to $8.5\%$ over the top-1 verifier. As shown in Fig.˜4, Weaver consistently outperforms naive ensemble averaging across both Oracle Top-5 Verifiers and Total Verifiers configurations for verifier ensembling, with improvements ranging from +2.4% to +10.1% across all datasets. The performance gains are particularly pronounced on GPQA Diamond (+10.1%) and MMLU Pro (+5.1%), demonstrating Weaver’s effectiveness in aggregating verifier signals through learned weights rather than simple averaging. However, gains diminish as more models are added—reflecting the classic ensemble bias-variance tradeoff: initial improvements stem from variance reduction, while additional verifiers contribute redundant signal due to correlated biases on hard examples abebuchanan2024 . We compare alternative score calibration strategies beyond Weaver’s binary transformation in Appendix C.7, and find that the default binarization yields the strongest downstream selection performance. We also explore scaling the number of scores per verifier—via prompt tuning or temperature variation—in Appendix C.5. While this yields modest improvements, increasing verifier count remains the more effective strategy. That said, both methods are complementary and can be combined for further gains.
@@
Figure 5: Weaver Improves the Accuracy-Compute Performance Trade-Offs. Success rate ($\%$) as a function of total inference compute per query (generation and verification compute, log scaled) for different verification strategies. Each point represents a different number of candidate generations (from $2^{0}$ to $2^{7}$). Weaver achieves the highest accuracy while requiring more compute than Majority Voting but demonstrates continued scaling benefits, while Weaver Distilled maintains most of Weaver’s performance gains with 97.3% compute savings and substantial accuracy improvements over baseline methods.
@@
We study how performance scales in the total compute used for both verification and repeated generations. Figure 5 shows the relationship between inference-time compute and success rate for different generation-verification systems. For each method, we scale the number of generations exponentially from 1 to 100 and plot the required inference compute for generation and verification together versus the success rate. Note that Fig.˜5 differs from Fig.˜3, since Majority Voting requires 0 verification inference calls while Weaver requires 30+ calls for the weak verifiers. We find that Weaver achieves the highest maximum success rate; notably, majority voting plateaus at around $2^{2}$ to $2^{3}$ ExaFLOPs per query while Weaver continues scaling until 512 ExaFLOPs. However, the additional compute required for Weaver can be prohibitive. We explore how to reduce this computational burden while retaining Weaver’s performance in the next section.
@@ 6 Weaver Distillation: Improving Verification Efficiency at Inference
Figure 6 shows the performance of Weaver on the Llama-70B generations against the cross-encoder on GPQA Diamond. Across tasks, we find that the distilled cross-encoder is able to capture 98.2% of the performance of Weaver. When running Weaver with all the verifiers, it costs 35.35 exaFLOPs for each query’s set of 100 samples. Running a 400M cross-encoder costs 1.01 exaFLOPs for evaluating 100 samples and reduces compute cost by more than three orders of magnitude, saving 99.97% of the FLOPs originally required for running the 70B verifiers. We also outperform majority voting by 23.2% while only incurring a 0.57% increased inference cost over only generating the responses. We see similar results for additional datasets in Figure 22 (Sec.˜C.6).
@@ B.1 Weak Supervision Model
[…]
@@
Eq.˜6 requires evaluating the full conditional likelihood:
@@
which is a joint distribution over $m$ binary random variables. Since each verifier $S_{k}\in\{0,1\}$ is binary, then there are $2^{m}$ possible verifier output configurations for $S\in\{0,1\}^{m}$. This results in $2^{m}-1$ free parameters per class label to construct the distribution $\Pr(S=\bar{s}_{ij}\mid y_{ij})$.
@@
Then, Eq.˜6 simplifies to a Naive Bayes-style estimator:
@@ B.2 Adapting Weak Supervision to the Verification Setting
Given a set of verifiers, we elaborate on the design choices behind the weak supervision model described in Sec.˜4. In particular, we describe challenges around normalization, binarization and filtering out low-quality verifiers. In Section B.3, we describe Weaver’s approach to normalizing, binarizing, and filtering verifiers.
@@ B.2.2 Binarization
The weak supervision algorithm described in the prior section requires binary verifier outputs. This is naturally suited for judge-style verifiers—such as language models prompted to answer yes/no questions—which output discrete 0,1{0,1} labels. However, many verifiers, especially reward models, emit continuous scores and often vary in scale and calibration. This raises a key design question: should we input these scores to the weak supervision model as-is, or should we binarize them?
Using continuous scores retains fine-grained information about the confidence of each verifier. This can improve ranking-based performance metrics such as AUC and may allow the weak supervision model to better resolve disagreements among verifiers. However, it introduces challenges when combining signals across verifiers with inconsistent calibration or scale: a score of 0.8 may have different meanings for different verifiers. As seen in Figure Fig.˜7 different verifiers exhibit different score distributions even when evaluating the same set of responses. Some verifiers are sharply bimodal, others skew heavily toward low or high scores, and some produce nearly flat or noisy distributions.
@@
Figure 8: Same conventions as Fig.˜9 where the performance reported is the area under the curve of the ROC.